OpenStudy (anonymous):

Given points A (1,1) and B (5,5), find all possible coordinates of point C that will make triangle ABC an equilateral.

OpenStudy (anonymous):

So far, I tried finding the distance between point A & B which I found out is radical 32... and then I don't know what to do after it

OpenStudy (anonymous):

I tried doing something like \[\sqrt{(x-1)^{2}+(y-1)^{2}} = \sqrt{32}\] \[\sqrt{(x-5)^{2}+(y-5)^{2}} = \sqrt{32}\] and then \[\sqrt{(x-1)^{2}+(y-1)^{2}} =\sqrt{(x-5)^{2}+(y-5)^{2}} \]

OpenStudy (dumbcow):

good start Now square both sides Then expand and add like terms solve for y in terms of x

OpenStudy (anonymous):

If done correctly, I got y=-x+6...? Is this correct?

OpenStudy (dumbcow):

yes it is, however i think i was wrong this will give you coordinates such that AC = BC but they won't equal sqrt(32) thus giving a isosceles triangle not an equilateral triangle hold on i think have the correct solution instead of setting them equal to each other you have to use substitution to solve system of 2 equations.

OpenStudy (curry):

i need your help dumb cow to check joe math's work

OpenStudy (anonymous):

What are the 2 equations?

OpenStudy (dumbcow):

Solve 1st equation for y: \[(x-1)^{2} + (y-1)^{2} = 32\] \[y = \pm \sqrt{32-(x-1)^{2}} +1\] substitute in for y in 2nd equation: \[(x-5)^{2}+(\sqrt{32-(x-1)^{2}}-4)^{2} = 32\] simplify \[\rightarrow \sqrt{32-(x-1)^{2}} = 5-x\] \[\rightarrow x^{2} - 6x -3 =0\] Use quadratic formula \[x = 3\pm2\sqrt{3}\]

OpenStudy (anonymous):

Sorry I got lost when you simplified \[\sqrt{32-(x-1)^{2}} \] How did it become 5-x?